You have a 77 test average in Math. Your mom wants the average up to an 80. You've taken 3 tests and you only have one test left. What do you need to get on that last test to raise your average to an 80?

Answer by checkley79(3341)

(3*77+X)/4=80
231+X=4*80
231+X=320
X=320-231
X=89 IS NEEDED ON THE 4 th. TEST TO AVERAGE 80.